OpenStudy (anonymous):

Angelica's service club is selling flowers for a fundraiser. Each flower costs them $0.50 to purchase. They set the sale price of each flower at the purchase price plus 40%. Write and solve an equation to determine how many flowers they will need to sell to make a profit of $500. Use complete sentences to explain what each number and variable in your equation represents. Write your solution in the context of the original problem.

OpenStudy (anonymous):

@PinkSapphire

OpenStudy (pinksapphire):

0.50x+.40=500

OpenStudy (pinksapphire):

To make 40% into a decimal you have to move the decimal point 2 places to the right.

OpenStudy (anonymous):

How do I know what the variable represents?

OpenStudy (pinksapphire):

x=the number of flowers.

OpenStudy (anonymous):

It asked "Use complete sentences to explain what each number and variable in your equation represents"...how do i figure that out?

OpenStudy (pinksapphire):

0.50 represents the price of each flower, 0.40 represents the sale price and 500 represents the amount of money Angelica wants to make.

OpenStudy (anonymous):

Oh, okay. Do you mind helping me with 3 more?

OpenStudy (pinksapphire):

Sure.

OpenStudy (anonymous):

Angelo took a group of people to the movies and spent $80 on tickets. Adult tickets cost $10 and child tickets cost $6. The equation to represent his cost is 10A + 6C = $80. If there were 5 children in the group, how many adult tickets did he purchase? Show your work and use a complete sentence to explain your solution in the context of the original problem.

OpenStudy (anonymous):

@PinkSapphire

OpenStudy (pinksapphire):

10A+6(5)=80 Plug in 5 for C. 10A+30=80 Subtract 30 on both sides. -30 -30 10A=50 10A/10=50/10 Divide by 10 on both sides to get A alone. A=5

OpenStudy (anonymous):

Oh, That was easier than I thought. Now, for the next two, I have tried SO many time but i cant get them right!!

OpenStudy (pinksapphire):

Angelo purchased 5 adult tickets.

OpenStudy (anonymous):

2. Max was solving the equation below and isn’t sure if his answer is correct. Explain to Max how he can check his answer and then help him identify any errors he made. Provide the correct solution in your explanation. 3x+6 = 7x-1 8 6 24 (3x+6) = 24(7x-1) 1 8 1 6 3(3x+6) + 4(7x-1) 9x+18 = 28x-1 9x-9x+18 = 28x-9x-1 18 = 19x-1 18+1 = 19x-1+1 19 = 19x 1 = x

OpenStudy (pinksapphire):

What do the 8 and 6 represent below the equation?

OpenStudy (anonymous):

Sorry, It is supposed to be fraction but i guess it didnt come up

OpenStudy (anonymous):

and both 24 is supposed to be over 1

OpenStudy (pinksapphire):

Okay :)

OpenStudy (pinksapphire):

He made a mistake of multiplying 3/8 to get 24/1 which is not right.

OpenStudy (anonymous):

where do you see that he multiplied 3/8 ?

OpenStudy (pinksapphire):

It doesn't really show it but 3*8=24

OpenStudy (anonymous):

Oh, Okay. And the last one: A quality control inspector is in charge of inspecting drill bits that are produced on an assembly line. The diameter of a drill bit can be modeled by the equation |x − 0.525| = 0.01. Show your work in solving the equation to determine the minimum and maximum diameters of the drill bits.

OpenStudy (pinksapphire):

I got x=0.535 and x=0.515

OpenStudy (pinksapphire):

This is how you do it: https://www.youtube.com/watch?v=NxbLXwXXp7Y

OpenStudy (anonymous):

Don't you set the answer to a negative a positive?

OpenStudy (pinksapphire):

No, just the equations.

OpenStudy (anonymous):

I attached my work from these problems..can you see if number four, the problem we just did, is correct?

OpenStudy (pinksapphire):

Sorry, I can't download it.